OpenStudy (anonymous):

One day, a person went to horse racing area, Instead of counting the number of human and horses, he instead counted 74 heads and 196 legs. Yet he knew the number of humans and horses there. How did he do it, and how many humans and horses are there?

OpenStudy (anonymous):

plzz help?

OpenStudy (anonymous):

wow this is an oldie!!

OpenStudy (dmezzullo):

well there is 4 legs per horse so 196 divided by 4 and heads meaning horse head or human?

OpenStudy (anonymous):

you can put \(x\) as the number of horses, and say \(y\) as the number of humans together they would have \(x+y=74\) heads and \(4x+2y=196\) legs

OpenStudy (anonymous):

your job is therefore to solve \[x+y=74\]\[4x+2y=196\] you have a choice of methods do you know how to solve this?